OpenStudy (anonymous):

State the vertex, direction of opening, axis of symmetry, y intercept, and maximum or minimum value for each of the following:

OpenStudy (anonymous):

\[y=(x+4)^{2}+1\]

OpenStudy (campbell_st):

expand the right hand side \[y = x^2 +8x + 17\] the coefficient of x^2 determines concavity.... +ve concave up -ve concave down the parabola is concave up x = -b/2a gives the line of symmetry of the parabola.... (-b/2a comes from the general quadratic formula) so x = -8/(2 x 1) the line of symmetry is x = -4 as the parabola is concave up it will have a minimum... The vertex is on the line x = -4 so substitute into the equation to find the coordinates of the vertex... (-4, 1) the y value from the above is the minimum value... to find the y- intercept, let x = 0 and solve.... intercept y = 17